5. Problems and Applications Q5 The nation of Textilia does not allow imports of clothing. In its equilibrium without trade, a T-shirt costs $20, and the equilibrium quantity is 3 million T-shirts. One day, after reading Adam Smiths The Wealth of Nations while on vacation, the president decides to open the Textilian market to international trade. The market price of a T-shirt falls to the world price of $16. The number of T-shirts consumed in Textilia rises to 4 million, while the number of T-shirts produced declines to 1 million. Use the blue line (circle symbol) to graph the domestic demand for T-shirts in Textilia. Then use the orange line (square symbol) to graph the domestic supply of T-shirts in Textilia. Next, use the black point (plus symbol) to indicate the domestic equilibrium price and quantity before trade Finally, use the grey line (star symbol) to indicate the world price. Note: Assume the domestic demand and domestic supply curves are linear. 40 36 Domestic Demand 32 28 Domestic Supply 24 20 16 Domestic Equilibrium 12 World Price Quantity of T-shirts (millions)Complete the following table by calculating the consumer surplus, producer surplus, and total surplus before and after trade. (Hint: Recall that the area of a triangle is x base x height.) Before Trade After Trade (Millions of Dollars) (Millions of Dollars) Consumer Surplus Producer Surplus Total Surplus ,producer surplus As a result of opening up to trade, consumer surplus surplus decreases ,and total
Answer #1

Graphs have been drawn correctly:

Before trade:

CS = (32 - 20) x 3/2 = $ 18 mn (area below demand curve and above price level)

PS = (20 - 14) x 3/2 = $ 9 mn (area above the supply curve and below price level)

TS = PS + CS = $ 27 mn

After trade :

CS = (32 - 16) x 4/2 = $ 32 mn

PS = (16 - 14) x 1/2 = $ 1 mn

TS = $ 33 mn

As a result of opening up to trade, consumer surplus increases, producer surplus decreases, and total surplus increases.